find y=log(x) if your domain is 0,1,2,3,4,5
Answer by Nate(3500)

y = log(0) = undefined
y = log(1) = 0
y = log(2) = 0.3010
y = log(3) = 0.4771
y = log(4) = 0.6021
y = log(5) = 0.6990